A man cycles to his destination with the wind at 8 ms-1


and returns home against


the wind at 4 ms-1


. What is his average speed? (Hint: Let the distance between home

and work


1
Expert's answer
2021-12-06T09:44:56-0500

When the object traveled two equal distances with different speed its average speed is given by

"v_{\rm ave}=\frac{2v_1v_2}{v_1+v_2}"

"v_{\rm ave}=\rm\frac{2*8\: m/s*4\: m/s}{8\: m/s+4\: m/s}=5.3\: m/s"